OpenStudy (anonymous):

solve the proportion for x. (x+4)/(2)=(4)/(x). What is the larger solution? What is the smaller solution?

OpenStudy (mertsj):

\[\frac{x+4}{2}=\frac{4}{x}\] \[x(x+4)=8\] \[x^2+4x=8\] \[x^2+4x-8=0\] \[x=\frac{-4\pm \sqrt{4^{2}-4(1)(-8)}}{2(1)}=\frac{-4\pm \sqrt{48}}{2}\]

OpenStudy (mertsj):

\[x=\frac{-4\pm 4\sqrt{3}}{2}=-2\pm2\sqrt{3}\]

OpenStudy (mertsj):

The larger root is \[-2+2\sqrt{3}\]

OpenStudy (mertsj):

The smaller root is \[-2-\sqrt{3}\]